34) - a) A=/0 3 il 14 1-1) 27 -5 Solve X=AX General solution) b) A=/6 6 6) Solve X=AX 11 -5 al (General Solution)
Answer #1

1 (a)A= 4 1 2 7 -5 -1 TO find the eigen values solve det (A– AI )= 0 |0-2 3 -1 =0=2 +42² – 122 – 80= 0=4 = 4, , = -4+2i, 2, =For 2 = -4+2i : (1 0 (1+2i)/5 (4-2i X = 0= 0 1 -(1+2i)/5 X = 0. -1- 2i) (4-(-4+2i)I)X =0→ 4 5- 2i -1 2 x +(1+2i)x, / 5= 0 [x,